# Dissecting the SWE-Bench Leaderboards: Profiling Submitters and Architectures of LLM- and Agent-Based Repair Systems

### v1

@@ Abstract.
The rapid progress in Automated Program Repair (APR) has been driven by advances in AI, particularly large language models (LLMs) and agent-based systems. SWE-Bench is a recent benchmark designed to evaluate LLM-based repair systems using real issues and pull requests mined from 12 popular open-source Python repositories. Its public leaderboards—SWE-Bench Lite and SWE-Bench Verified—have become central platforms for tracking progress and comparing solutions. However, because the submission process does not require detailed documentation, the architectural design and origin of many solutions remain unclear. In this paper, we present the first comprehensive study of all submissions to the SWE-Bench Lite (68 entries) and Verified (79 entries) leaderboards, analyzing 67 unique approaches across dimensions such as submitter type, product availability, LLM usage, and system architecture. Our findings reveal the dominance of proprietary LLMs (especially Claude 3.5/3.7), the presence of both agentic and non-agentic designs, and a contributor base spanning from individual developers to large tech companies.
@@ 1. Introduction
The field of Automated Program Repair (APR) has progressed rapidly over the last decade. Early approaches relied on techniques such as search-based methods—for example, GenProg (LeGoues2011genprog, )—and constraint-based methods, including SemFix (nguyen2013Semfix, ) and Nopol (Xuan2017nopol, ). More recently, advances in Artificial Intelligence—particularly in deep learning and transformer architectures—have significantly influenced the field. Many current state-of-the-art techniques (ruan2024specrovercodeintentextraction, ; ma2025alibabalingmaagent, ; bouzenia2024RepairAgent, ) are based on agents powered by large language models (LLMs), which can autonomously handle the entire repair process, from understanding the issue to validating the generated patch.
@@
This progress has been significantly facilitated by the creation of bug benchmarks, which have allowed researchers to focus more on developing repair techniques rather than on how to evaluate them. They also serve as a common ground for evaluating and comparing the performance of APR techniques. One of the most widely adopted benchmarks is Defects4J (Just2014Defects4J, ), which contains over 800 real bugs mined from open-source Java libraries. It also provides a framework for reproducing bugs, visualizing ground truth patches, and other essential functionalities. Defects4J has been widely used to evaluate a broad range of techniques, from traditional search-based approaches (martinez2017automatic, ; Liu2020:ontheefficiency, ) to emerging methods, including deep learning-based techniques (deeprepair2019, ; Chen2021SequenceR, ), transformer models (Ye2022RewardRepair, ; Smytzek2024FixKit, ; Lutellier2020Coconut, ; Nan2021Cure, ), early explorations with large language models (LLMs) (Jiang2023Impact, ; Xia2024ChatRepair, ), and, more recently, agentic APR systems (bouzenia2024RepairAgent, ; ye2025adversarial, ).
Recently, new benchmarks have emerged to address evolving scenarios—most notably SWE-Bench (jimenez2024SWEBenchLLMs, ), which includes 2,294 instances mined from pull requests and issues across 12 widely used Python repositories.
@@
SWE-Bench differs from Defects4J in several important ways. First, the instances in SWE-Bench are written in Python, whereas Defects4J focuses on Java. Second, each instance in SWE-Bench does not come with test cases or scripts that expose the bug and validate the fix. In contrast, Defects4J includes such tests (these test were mined from the commits that also introduce the fix). Third, SWE-Bench provides a framework to run experiments and compute metrics such as precision (percentage of resolved issues) and project-level repair statistics, similar to what projects like RepairThemAll (Durieux:2019:RepairThemAll, ) have built on top of Defects4J. Finally, the creators of SWE-Bench actively maintain public leaderboards where evaluation results are published and compared. Researchers and developers can submit their results via pull requests, including output generated by the evaluation framework and accompanying execution logs.
@@
The initial results published on the SWE-Bench leaderboard correspond to experiments conducted by its authors to assess the repair capabilities of large language models, specifically, a fine-tuned version of CodeLlama (roziere2024codellama, ). Subsequently, the same authors introduced and evaluated an approach based on an Agent backed by a LLM (SWE-Agent). A key feature of SWE-Bench is the inclusion of issue descriptions, which enables the evaluation of LLM-based and agentic systems in reasoning over natural language and source code, and generating patches. Since its release, SWE-Bench has been adopted by researchers to benchmark their repair approaches and the experimental results included in their articles (e.g., (Zhang2024AutoCoderRover, ; ma2025alibabalingmaagent, ; xia2024agentlessdemystifyingllmbasedsoftware, ; ruan2024specrovercodeintentextraction, ; li2025patchpilot, )). Moreover, many of them have submitted their results to the leaderboard.
@@
The SWE-Bench leaderboard enables the research community —and the public at large— to regularly track the progress of intelligent systems (e.g., those based on LLMs or agents) in addressing software issues. While this openness to submissions plays a crucial role in fostering innovation, it introduces a limitation from a research perspective: submissions are not required to disclose how the reported results were obtained. For example, although some entries are linked to academic publications such as AutoCoderRover (Zhang2024AutoCoderRover, ) or Agentless (xia2024agentlessdemystifyingllmbasedsoftware, ), many others are not. For instance, there are submissions from Google and Amazon. Rather than viewing this as a weakness of the leaderboard, we argue that it enables broader participation by organizations beyond academia, such as startups and tech companies, who may wish to showcase the capabilities of their technologies while protecting their intellectual property. This lack of publicly available details has motivated our study. In this paper, we aim to analyze and categorize the solutions submitted to the SWE-Bench leaderboard, in order to better understand the architectural choices and features that are driving progress in automated program repair.
In this paper, we present the first in-depth study of the SWE-Bench leaderboard. Our primary goal is to understand who is submitting to the leaderboard and what types of solutions are being submitted. We focus on the two main leaderboards hosted on the SWE-Bench platform: (1) SWE-Bench Lite, which consists of 300 instances selected from the full benchmark; and (2) SWE-Bench Verified, which includes 500 instances curated and verified by OpenAI (chowdhury2024swebenchverified, ).
@@
To conduct our study, we examined each entry in the SWE-Bench leaderboards and collected the available metadata, including the submission pull request, submitter link, and precision score (Resolved). We then searched both academic publications and grey literature (e.g., blog posts, LinkedIn posts) to gather additional information about each submission. This includes the nature of the submitter (e.g., company or academic institution), any software artifact related to the submission, the LLMs used, the openness of the solution, and relevant architectural details of the solution. To describe the high-level architecture of each submitted approach, we focus on three dimensions: (1) the repair workflow authoring, (2) the autonomy over the execution path, and (3) the number of LLM-based Agents, if the solution is based on Agents (for agent-based solutions).
@@
In total, we analyzed 68 entries from SWE-Bench Lite, 79 entries from SWE-Bench Verified, and 67 unique approaches. [^1] An approach can be evaluated on both leaderboards, and some approaches have multiple entries, e.g., using different LLMs. Our results show that, while there is a diversity of submitter types (e.g., academia, industry, and collaborations), the majority of submissions come from industry , particularly from small companies and large publicly traded corporations such as Amazon, IBM, and Google. We also identified submissions made by individual developers, highlighting how access to powerful language and code models —whether commercial (e.g., OpenAI) or open-source (e.g., Meta)— is enabling the creation of sophisticated repair systems.
We also found that the majority of submissions achieving state-of-the-art results rely on proprietary LLMs —most notably, Claude 3.5/3.7 Sonnet, which currently dominates the leaderboard. In terms of system architecture, submissions are highly diverse: they range from single-LLM solutions without agentic capabilities to complex multi-agent systems with emergent workflows. Notably, each of these architectural categories includes at least one submission that achieves near state-of-the-art performance. To the best of our knowledge, this paper presents the first comprehensive characterization of all submitters and solutions published on the SWE-Bench leaderboards.
The research questions that guide our study are as follows:
@@
The paper continues as follows: Section 2 presents the methodology of our study. Section 3 presents the results. Section 4 presents a discussion of the results. Section 5 presents the related work. Section 6 concludes the paper.
@@ 2. Methodology
In this section we describe the methods we follow to respond to the research questions. In Section 2.1 we present the method to inspect the SWE-Bench dashboards and to respond to RQ 1. Section 2.2 outlines the method used to analyze the submitted approaches in relation to RQ2 and RQ3. Finally, Section 2.3 details the data collected to support our analysis and answer the research questions.
@@ 2.1.1. Data Collection Procedure
To collect the data, we inspect the SWE-Bench Lite and SWE-Bench Verified leaderboards as follows. First, we visit each leaderboard page and retrieve from each entry (i.e., each submission): Submission name, Resolved rate (%), and Site (URL). Then, for each entry, we conduct the following steps. a) We visit the listed site to gather valuable information about the approach (such as its architecture and design decisions) and to find any references related to experiments conducted on SWE-Bench. b) While browsing a page, we manually navigate the entire site and follow any external links that may be useful for our goal (e.g., link to a blog or a paper). c) We supplement these data with a dedicated Google search to capture additional pages that are not accessible through the website linked to the leaderboard. The Google query we created is "<Name_Entry> + SWE-Bench"; in "<Name_Entry>", we remove the date and the model’s details. For example, given the entry "Gru(2024-12-08)", the name used in the Google search is "GRU". d) We inspect the pages obtained with the web search, and store those that describe the approach or the experiment conducted on SWE-Bench. e) We also inspect the data submitted for each entry, notably the README.md [^2] Example of README.md file included in a commit to SWE-Bench: evaluation/lite/20240509_amazon-q-developer-agent-20240430-dev/README.md and metadata.yaml files, available in the SWE-Bench’s repository [^3] https://github.com/SWE-bench/experiments/tree/main/evaluation/lite dedicated to collect the row information from the submissions. f) We also visit the LinkedIn of the submitter, and eventually some of its members, in order to detect posts that discuss the experiment. In addition to collecting the submitter information for each entry, we also identify the underlying approach associated with it. For example, the entries "SWE-agent + GPT-4o (2024-05-13)" and "SWE-agent + GPT-4 (1106)" originate from the same research group (Princeton University) and are based on the same approach, SWE-agent, but differ in the LLM used (GPT-4o and GPT-4, respectively).
@@ 2.1.2. Description of data to collect
    Academia: The submission is made by members affiliated with academic institutions and is accompanied by a research article whose authors are all from academia. We distinguish among single Academia institutions and Collaboration Academia, when different academic institutions collaborate in a submissions.
@@
      Unknown: Companies that we were unable to subclassify due to lack of information.
@@
  • The Product Purpose: describes what the product does. We have identified ten different purpose, and an additional code is introduced for entries without product. a) Agent Framework, b) Coding Assistant, c) Development Assistant, d) Development Framework, e) Development Platform, f) Issue Resolution, g) Other-Inference Service, h) Platform Code Optimization/Improvement, i) Platform Code Representation, and j) Platform Integration.
  • The Product form: Describes how the product is delivered or used. We have identified five different forms, and an additional code is introduced for unknown cases. a) Cloud platform, b) Command-line tool, c) Github plugin, d) IDE plugin, and e) Library.
@@
Second, we record the availability of the product. Our deductive coding on both leaderboards resulted in four codes representing different levels of availability:
@@
    Publicly Available Product (PAP): A product based on the submitted solution is publicly accessible, either commercially (paid) or for free. This includes IDE plugins and cloud-based applications (e.g., cloud-based IDEs). The product must be accessible online, for example, as an IDE plug-in available in plug-in/extensions marketplaces (such as those for Visual Studio [^7] https://marketplace.visualstudio.com/vscode or IntelliJ [^8] https://plugins.jetbrains.com/ nor a deployed solution such as online IDE. Importantly, this category does not refer to the availability of source code; for instance, an IntelliJ plugin may be either open- or closed-source.
@@
Note that in this analysis we focus on products associated to each entry -if any- rather that on the code used in the SWE-Bench experiment (which could be eventually the same).
@@
2.2. Inspection of approaches listed in SWE-Bench leaderboard
In the previous step, we collected artifacts related to each leaderboard entry, which include -if available- published papers, ArXiv documents, Blog post, LinkedIn post, documentation on Github repository (Readme.me) or information contained in the commit done for the submission. We now analyze these artifacts in order to extract information that describes the approaches associated to the submissions.
In particular the information that we aim at capturing the architecture of the approaches. For that, we focus on different dimensions.
2.2.1. Number of Agents
The first dimension is Number of Agents. A fast observation of the leaderboard entries reveals, at least by the name, that most of the solutions are based on Agents (e.g., SWE-Agent). Nevertheless, we define the following categories: a) No Agent. This includes repair systems that just prompts LLMs e.g., using Chain of Thoughts. b) Single Agent. c) Multiple Agents.

### v2

@@ Abstract.
The rapid progress in Automated Program Repair (APR) has been driven by advances in AI, particularly large language models (LLMs) and agent-based systems. SWE-Bench is a recent benchmark designed to evaluate LLM-based repair systems using real issues and pull requests mined from 12 popular open-source Python repositories. Its public leaderboards—SWE-Bench Lite and SWE-Bench Verified—have become central platforms for tracking progress and comparing solutions. However, because the submission process does not require detailed documentation, the architectural design and origin of many solutions remain unclear. In this paper, we present the first comprehensive study of all submissions to the SWE-Bench Lite (79 entries) and Verified (99 entries) leaderboards, analyzing 80 unique approaches across dimensions such as submitter type, product availability, LLM usage, and system architecture. Our findings reveal the dominance of proprietary LLMs (especially Claude 3.5), the presence of both agentic and non-agentic designs, and a contributor base spanning from individual developers to large tech companies.
@@ 1. Introduction
The field of Automated Program Repair (APR) has progressed rapidly over the last decade. Early approaches relied on techniques such as search-based methods —for example, GenProg (LeGoues2011genprog, )— and constraint-based methods, including SemFix (nguyen2013Semfix, ) and Nopol (Xuan2017nopol, ). More recently, advances in Artificial Intelligence (AI) —particularly in deep learning and transformer architectures— have significantly influenced the field. Many current state-of-the-art techniques (ruan2024specrovercodeintentextraction, ; ma2025alibabalingmaagent, ; bouzenia2024RepairAgent, ) are based on agents powered by large language models (LLMs), which can autonomously handle the entire repair process, from understanding the issue to validating the generated patch.
@@
This progress has been significantly facilitated by the creation of bug benchmarks, which have allowed researchers to focus more on developing repair techniques rather than on how to evaluate them. They also serve as a common ground for evaluating and comparing the performance of APR techniques. One of the most widely adopted benchmarks is Defects4J (Just2014Defects4J, ), which contains over 800 real bugs mined from open-source Java libraries. It also provides a framework for reproducing bugs, visualizing ground truth patches, and other essential functionalities. Defects4J has been widely used to evaluate a wide range of techniques, from traditional search-based approaches (martinez2017automatic, ; Liu2020:ontheefficiency, ) to emerging methods, including deep learning-based techniques (deeprepair2019, ; Chen2021SequenceR, ), transformer models (Ye2022RewardRepair, ; Smytzek2024FixKit, ; Lutellier2020Coconut, ; Nan2021Cure, ), early explorations with LLMs (Jiang2023Impact, ; Xia2024ChatRepair, ), and, more recently, agentic APR systems (bouzenia2024RepairAgent, ; ye2025adversarial, ).
@@
Recently, new benchmarks have emerged to address evolving scenarios, most notably SWE-Bench (jimenez2024SWEBenchLLMs, ), which includes 2,294 instances mined from pull requests and issues across 12 widely used Python repositories. SWE-Bench differs from Defects4J and similar benchmarks (e.g. (Madeiral2019bears, ; lin20217Quixbugs, )) in several important ways. First, each instance in SWE-Bench does not come with test cases or scripts that expose the bug and validate candidate fixes. In contrast, Defects4J provides such tests —mined from the same commits that introduce the fix or from subsequent ones. The SWE-Bench setup is more aligned with real-world scenarios that developers typically face, where test cases are not available when a bug is reported. Second, SWE-Bench provides a framework for running experiments and computing metrics such as precision (percentage of resolved issues) and project-level repair statistics, similar to what projects like RepairThemAll (Durieux:2019:RepairThemAll, ) have built on top of Defects4J and other benchmarks. Finally, the creators of SWE-Bench actively maintain public leaderboards where evaluation results are published and compared. Researchers and developers can submit their results via pull requests, including output generated by the evaluation framework and accompanying execution logs. Such leaderboards were uncommon in earlier APR research, which often lacked platforms for sharing and comparing results.
@@
The initial results published on the SWE-Bench leaderboard correspond to experiments conducted by its authors to assess the repair capabilities of large language models, specifically, a fine-tuned version of CodeLlama (roziere2024codellama, ). Subsequently, the same authors introduced and evaluated an approach based on an Agent backed by an LLM (SWE-Agent). A key feature of SWE-Bench is the inclusion of issue descriptions, extracted from the GitHub platform, which enables the evaluation of LLM-based and agentic systems in reasoning over both natural language and source code, and in generating patches. Since its release, SWE-Bench has been adopted by researchers to benchmark their repair approaches and the experimental results included in their articles (e.g., (Zhang2024AutoCoderRover, ; ma2025alibabalingmaagent, ; xia2024agentlessdemystifyingllmbasedsoftware, ; ruan2024specrovercodeintentextraction, ; li2025patchpilot, )). Moreover, many of them have submitted their results to the leaderboards.
@@
The SWE-Bench leaderboard enables the research community —and the public at large— to regularly track the progress of intelligent systems (e.g., those based on LLMs or agents) in addressing software issues. While this openness to submissions plays a crucial role in fostering innovation, it introduces a limitation from a research perspective: submissions are not required to disclose how the reported results were obtained. For example, although some entries are linked to academic publications such as AutoCoderRover (Zhang2024AutoCoderRover, ) or Agentless (xia2024agentlessdemystifyingllmbasedsoftware, ), many others (e.g., from IBM or Amazon) are not. Rather than viewing this as a weakness of the leaderboard, we argue that it enables broader participation by organizations beyond academia, such as startups and tech companies, who may wish to showcase the capabilities of their technologies while protecting their intellectual property. This lack of publicly available details has motivated our study. In this paper, we present the first in-depth study of the SWE-Bench leaderboards. We aim to analyze and categorize the solutions submitted to them, in order to learn who are the contributors to the benchmark and better understand the architectural choices and features that are driving progress in automated program repair. We focus on two leaderboards hosted on the SWE-Bench website: (1) SWE-Bench Lite, which consists of 300 instances selected from the full benchmark; and (2) SWE-Bench Verified, which includes 500 instances curated and verified by OpenAI (chowdhury2024swebenchverified, ).
@@
To carry out our study, we examine each entry in the SWE-Bench leaderboards and collect the available metadata, including the submission pull request, submitter link, and precision score (% Resolved). We then searched both academic publications and grey literature (e.g., blog posts, LinkedIn posts) to gather additional information about each submission. This includes the nature of the submitter (e.g., company or academic institution), any software artifact related to the submission, the LLMs used, the openness of the solution, and relevant architectural details of the solution. To describe the high-level architecture of each submitted approach, we focus on three dimensions: (1) the repair workflow authoring, (2) the autonomy over the execution path, and (3) the number of LLM-based Agents, if the solution is agent-based.
@@
In total, we analyze 79 entries from SWE-Bench Lite, 99 entries from SWE-Bench Verified, comprising 80 unique approaches. [^1] An approach can be evaluated on both SWE-Bench Lite and SWE-Bench Lite leaderboards, and some approaches have multiple entries, e.g., each using different LLMs. Our results show that, while there is a diversity of submitter types (e.g., academia, industry, and collaborations), the majority of submissions come from industry, particularly from small companies and large publicly traded corporations such as Amazon, IBM, and Google. We also identified submissions made by individual developers, highlighting how access to powerful language and code models —whether commercial (e.g. GTP 4 from OpenAI, Claude 4 from Anthropic) or open-source (e.g., Llama from Meta, Qwen from Alibaba)— is enabling the creation of sophisticated repair systems.
We also found that the majority of submissions achieving state-of-the-art results rely on proprietary LLMs —most notably, Claude 4 Sonnet. In terms of system architecture, submissions are highly diverse: they range from single-LLM solutions without agentic capabilities, to complex multi-agent systems with emergent workflows. Notably, no single architecture consistently achieves state-of-the-art performance. Rather, high-performing submissions follow a variety of architectural strategies, suggesting that multiple design paradigms can be effective. We also present a discussion of SWE-Bench that covers patch quality concerns, debates over agentic architectures, the openness of solutions, as well as limitations and potential saturation of the benchmark. To the best of our knowledge, this paper presents the first comprehensive characterization of all submitters and solutions published on the SWE-Bench leaderboards. In fact, the type of analysis of architectures and pipelines that we have carried out could be eventually applied to similar benchmarks.
The research questions that guide our research are:
@@
The paper continues as follows. Section 2 presents the methodology of our study. Section 3 presents the results. Section 4 presents a discussion of the results. Section 5 presents the threats to validity. Section 6 presents the related work. Section 7 concludes the paper.
@@ 2. Methodology
In this section, we describe the methods used to answer the research questions. In Section 2.1 we present the method to inspect the SWE-Bench dashboards and to respond to RQ 1. Sections 2.2 and 2.3 outline the methods used to analyze the submitted approaches in relation to RQ2 and RQ3. Finally, Section 2.4 details the data collected to support our analysis and answer the research questions.
@@ 2.1.1. Data Collection Procedure
To collect data, we inspect the SWE-Bench Lite and SWE-Bench Verified leaderboards as follows. First, we visit each leaderboard page and retrieve from each entry (that is, each submission): Submission name, Resolved rate (%),and Site (URL). Then, for each entry, we conduct the following steps. a) We visit the listed site to gather valuable information about the approach (such as its architecture and design decisions) and to find any references related to experiments conducted on SWE-Bench. b) While browsing a page, we manually navigate the entire site and follow any external links that may be useful for our goal (e.g., link to a blog or a paper). c) We supplement these data with a dedicated Google search to capture additional pages that are not accessible through the website linked to the leaderboard. The Google query we create is "<Name_Entry> + SWE-Bench"; in "<Name_Entry>", we remove the date and the model’s details. For example, given the entry "Gru(2024-12-08)", the name used in the Google search is "GRU". d) We inspect the pages obtained with the web search, and store those that describe the approach or the experiment conducted on SWE-Bench. e) We also inspect the data submitted for each entry, notably the README.md [^2] Example of README.md file included in a commit to SWE-Bench: evaluation/lite/20240509_amazon-q-developer-agent-20240430-dev/README.md and metadata.yaml files, available in the SWE-Bench’s repository [^3] https://github.com/SWE-bench/experiments/tree/main/evaluation/lite dedicated to collect the row information from the submissions. f) We also visit the LinkedIn of the submitter, and eventually some of its members, in order to detect posts that discuss the experiment. In addition to collecting the submitter information for each entry, we also identify the underlying approach associated with it. For example, the entries "SWE-agent + GPT-4o (2024-05-13)" and "SWE-agent + GPT-4 (1106)" originate from the same research group (Princeton University) and are based on the same approach, SWE-agent, but differ in the LLM used (GPT-4o and GPT-4, respectively). We collected information about the LLMs used in each submission from multiple sources: a) the submission name (e.g., GPT-4o and GPT-4, as in the entries mentioned above); b) the metadata.yaml file—if available—specifically the model field; c) the README.md file; d) associated scientific articles and grey literature (blog posts).
Availability and configuration of data has sometimes restricted the type of analysis. For instance, we initially planned an analysis of contributions per country, but at the end we discarded this option since it was not entirely clear in many leaderboard entries.
@@ 2.1.2. Description of data to collect
    Academia: The submission is made by members affiliated with academic institutions and is accompanied by a research article whose authors are all from academia. We distinguish among Single Academy institutions and Collaboration Academia, when different academic institutions collaborate in a submissions.
@@
      Unknown: Companies for which we were unable to subclassify due to lack of information.
@@
  • Product Purpose: describes what the product does. We identify ten different purposes: a) Agent Framework, b) Coding Assistant, c) Development Assistant, d) Development Framework, e) Development Platform, f) Issue Resolution, g) None, h) Other-Inference Service, i) Platform Code Optimization/Improvement, j) Platform Code Representation, k) Platform Integration, and l) Problem Solving. An additional code is introduced for products for which we cannot identify the form.
  • Product Form: Describes how the product is delivered or used. We identify five different forms: a) Cloud platform, b) Command-line tool, c) Github plugin, d) IDE plugin, and e) Library. An additional code is introduced for entries without form.
@@
Second, we record the availability of the product. Our deductive coding on both leaderboards results in four codes representing different levels of availability:
@@
    Publicly Available Product (PAP): A product based on the submitted solution is publicly accessible, either commercially (paid) or for free. This includes plugins for desktop IDEs (e.g., IntelliJ) and cloud-based applications (e.g., cloud-based IDEs). The product must be publicly accessible, for example, as an IDE plug-in available in plug-in/extensions marketplaces, such as those for Visual Studio [^7] https://marketplace.visualstudio.com/vscode or IntelliJ [^8] https://plugins.jetbrains.com/. Importantly, this category does not refer to the availability of source code; for instance, an IntelliJ plugin may be either open- or closed-source.
@@
Note that in this analysis, we focus on products associated to each entry -if any- rather than on the code used in the SWE-Bench experiment (which could be eventually the same).
@@
2.2. Architecture of approaches in SWE-Bench
In the previous step, we collected artifacts related to each leaderboard entry, which include -if available- published papers, ArXiv documents, Blog post, LinkedIn post, documentation on Github repository (Readme.me) or information contained in the commit done for the submission. We now analyze these artifacts in order to extract information that describes the approaches associated with the submissions.
In RQ2, the information that we aim at capturing is the architecture of the approaches. For that, we have elicited three different dimensions.